A study is conducted to investigate the impact of depression on a patient’s ability to survive cardiac disease. They evaluated 477 participants for depression and followed for 4 years. There were 374 patients with no depression (ND) and 69 of them died. Out of 103 patients with minor or major depression (D), 28 died. Construct a 95% confidence interval for the difference in proportion of deaths between depressed patients with cardiac disease and non-depressed patients with cardiac disease. (Hint: Round sample proportions and SE to four decimal places.)

A) (0.0392, 0.1356)

B) (-0.0072, 0.1818)

C) (0.0080, 0.1666)

D) (-0.0371, 0.2118)

E) (-0.0572, 0.2320)
Answer #1

Solution:

Given:

depressed patients with cardiac disease:

n1 = 103

x1=28

non-depressed patients with cardiac disease.

n2 = 374

x2 = 69

We have to find  a 95% confidence interval for the difference in proportion of deaths between depressed patients with cardiac disease and non-depressed patients with cardiac disease.

Formula:

(pi - p2 - E, pi-p2 + E)

where

P1 = în _ 11 28 = 0.2718 = 103

6-29 2 = = 374 = 0.1845

E = 2 x S Epipi

SEpi-Pa = Px (1 –p) – P2 x (1 – 12) ni n2

0.2718 x (1 – 0.2718) SEpi-p2 = V 0.1845 x (1 – 0.1845) 374 103

SEpi-p = 0.00192180 + 0.00040229

SEpi-p = 0.00232408

SEpi-p = 0.0482

We need to find zc value for c=95% confidence level.

Find Area = ( 1 + c ) / 2 = ( 1 + 0.95) /2 = 1.95 / 2 = 0.9750

Look in z table for Area = 0.9750 or its closest area and find z value.

0.1 V123 .00 .01 .02 .03 .04 .05 .06 0.0 ,5000 15040 5080 5120 .5160 5199 39 1.5398 .5438 5478 1.5517 .5557 .5596 .636 0.2 15

Area = 0.9750 corresponds to 1.9 and 0.06 , thus z critical value = 1.96

That is : Zc = 1.96

Thus

E = 2 x S Epipi

E = 1.96 x 0.0482

E = 0.0945

Thus

(pi - p2 - E, pi-p2 + E)

(0.2718 -0.1845 -0.0945,0.2718 -0.1845 +0.0945)

(-0.0072.0.1818)

Thus correct option is:

B) (-0.0072, 0.1818)